OpenStudy (anonymous):

for the data in each table, tell whether y varies directly with x. if it does write an equation for the direct variation x: -3 -1 2 5 y: 7.5 2.5 2 -12.5

OpenStudy (cwrw238):

7.5 / -3 = -2.5 2.5 / -1 = -2.5 2 / 2 = 1 -12.5 / 5 = -2.5 are yoy sure that x = 2 for y = 2 is correct?

myininaya (myininaya):

thats what i was wondering

OpenStudy (cwrw238):

if for instance for x = 2 y = -5 then the relationship would be y = -2.5 x

OpenStudy (anonymous):

oh thanks guys its x=2 y=-5

OpenStudy (cwrw238):

yw